OpenStudy (anonymous):

Uhhh, hate this helpp anyone? Solve the equation. 5d=17.1

OpenStudy (anonymous):

divide by 5

OpenStudy (anonymous):

5d=17.1 /5 /5 d =3.42

OpenStudy (ghazi):

\[\frac{ 5d }{ 5 }=\frac{ 17.1 }{ 5 }\]

OpenStudy (anonymous):

Thank you very much @surdawi